How do you solve: y/12 = (-8)^2 - l 21 l + (-4)
The 21 is absolute. Cannot find anything on how to perform this.
Thank you.

Answer by wuwei96815(245)

y/12 = (-8)^2 -[21] + (-4)
y/12 = 64 - 21 -4 = 39
y = 39 x 12
y = 468

Is that right?